# Approximate independence of permutation mixtures

### v1

@@ 3.3 A key inequality
is multilinear in $(r_{1},\ldots,r_{\ell})$. We invoke the following deep result due to S. Banach.
Lemma 3.2 ([Ban38]).
@@ 5.2 Application to a Gaussian sequence model
The contiguity relation $\{\mathbb{P}_{n}\}\triangleleft\{\mathbb{Q}_{n}\}$ in Theorem 1.2, or more specifically the mutual information relation in Corollary 1.4, can be applied in statistical problems. We review a classical example of a Gaussian sequence model as follows. Let the observation vector be $X\sim{\mathcal{N}}(\theta,I_{n})$, with an unknown mean vector $\theta\in\mathbb{R}^{n}$ lying in the following $\ell^{p}$ ball $\Theta_{p}(R)$:
@@ Acknowledgments.
We are grateful to Terence Tao for suggesting the use of the saddle point method in the proof of Theorem 3.3, to Cun-Hui Zhang for introducing us the question of the Bayes risk under a permutation prior in Lemma 5.3 and the reference [Zha12], and to Yunzi Ding and Cheng Mao for discussions about an earlier version of this project. JNW is partially supported by a Sloan Research Fellowship and NSF grant DMS-2210583.
@@ Appendix C Deferred proofs
It is clear that
@@
$={\sf KL}(P_{X}\|Q_{X}),$
where $P_{X},Q_{X}$ are the marginal distributions of the random vector $X$ under the permutation prior and the i.i.d. prior of $\vartheta$, respectively, and we have used $P_{X_{1}}=Q_{X_{1}}$ in the last step. So clearly $nI(\vartheta_{1};X_{1})-I(\vartheta;X)\geq 0$, and
@@
$\leq\min\left\{10\sum_{\ell=2}^{n}I_{\chi^{2}}(\vartheta_{1};X_{1})^{\ell},I_{\chi^{2}}(\vartheta_{1};X_{1})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})(\{P_{\theta_{1}},\dots,P_{\theta_{n}}\}))\right\},$
where we invoke Theorem 1.2 for ${\mathcal{P}}=\left\{P_{\theta_{1}},\dots,P_{\theta_{n}}\right\}$ and replace the channel capacity ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ by a tighter quantity $I_{\chi^{2}}(\vartheta_{1};X_{1})$, which is an intermediate step in the proof of Theorem 1.2.
@@
For the difference in the mutual information, Corollary 1.4 states that

### v2

@@ 3.3 A key inequality
is multilinear in $(r_{1},\ldots,r_{\ell})$. We invoke the following deep result due to S. Banach [Ban38].
Lemma 3.2.
@@ 5.2 Application to a Gaussian sequence model
The contiguity relation $\{\mathbb{P}_{n}\}\triangleleft\{\mathbb{Q}_{n}\}$ in Theorem 1.2, or more specifically the mutual information relation (1) in Corollary 1.4, can be applied in statistical problems. We review a classical example of a Gaussian sequence model as follows. Let the observation vector be $X\sim{\mathcal{N}}(\theta,I_{n})$, with an unknown mean vector $\theta\in\mathbb{R}^{n}$ lying in the following $\ell^{p}$ ball $\Theta_{p}(R)$:
@@ Acknowledgments.
We are grateful to Terence Tao for suggesting the use of the saddle point method in the proof of Theorem 3.3, to Cun-Hui Zhang for introducing us the question of the Bayes risk under a permutation prior in Lemma 5.3 and the reference [Zha12], to Yury Polyanskiy for pointing us to the reference [TP23], and to Yunzi Ding and Cheng Mao for discussions about an earlier version of this project. JNW is partially supported by a Sloan Research Fellowship and NSF grant DMS-2210583.
@@ Appendix C Deferred proofs
For the first inequality (1), it is clear that
@@
=$\overset{\rm(a)}{=}{\sf KL}(P_{X}\|Q_{X}),$
where $P_{X},Q_{X}$ are the marginal distributions of the random vector $X$ under the permutation prior and the i.i.d. prior of $\vartheta$, respectively, and (a) uses $P_{X_{1}}=Q_{X_{1}}$. So $nI(\vartheta_{1};X_{1})-I(\vartheta;X)\geq 0$, and
@@
$\leq\min\left\{10\sum_{\ell=2}^{n}I_{\chi^{2}}(\vartheta_{1};X_{1})^{\ell},I_{\chi^{2}}(\vartheta_{1};X_{1})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))\right\},$
where we use Theorem 1.2 and replace the channel capacity ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ by a tighter quantity $I_{\chi^{2}}(\vartheta_{1};X_{1})$, which is an intermediate step in the proof of Theorem 1.2.
For the second inequality (8), let $Q_{1},\dots,Q_{m}$ be an $\varepsilon$-cover of $\mathrm{conv}({\mathcal{P}})$ under the KL divergence, with a fixed $\varepsilon>0$ and $m=N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)$. By the standard entropic upper bound of the mutual information (cf. [YB99] or [PW24, Theorem 32.4]), one has
$I(\theta;X)$ $\leq\mathbb{E}_{\theta}\left[\min_{i\in[m]}{\sf KL}(P_{X|\theta}\|Q_{i}^{\otimes n})\right]+\log m$
=$\overset{\rm(b)}{=}\mathbb{E}_{\theta}\left[{\sf KL}(P_{X|\theta}\|Q_{X|\theta})+\min_{i\in[m]}\mathbb{E}_{P_{X|\theta}}\left[\log\frac{\mathrm{d}Q_{X|\theta}}{\mathrm{d}Q_{i}^{\otimes n}}\right]\right]+\log m$
=$\overset{\rm(c)}{=}\mathbb{E}_{\theta}\left[{\sf KL}(P_{X|\theta}\|Q_{X|\theta})+\min_{i\in[m]}{\sf KL}(Q_{X|\theta}\|Q_{i}^{\otimes n})\right]+\log m$
$\overset{\rm(d)}{\leq}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))+n\varepsilon^{2}+\log m,$
where in (b) $Q_{X|\theta}$ denotes the i.i.d. distribution $\left(\frac{1}{n}\sum_{i=1}^{n}P_{\theta_{i}}\right)^{\otimes n}$, (c) uses the i.i.d. structures for both $Q_{X|\theta}$ and $Q_{i}^{\otimes n}$, and notes that the marginal distributions $P_{X_{1}|\theta}$ and $Q_{X_{1}|\theta}$ are the same, and (d) follows from Theorem 1.2 and the definition of the KL covering applied to $\frac{1}{n}\sum_{i=1}^{n}P_{\theta_{i}}\in\mathrm{conv}({\mathcal{P}})$. Taking the infimum over $\varepsilon>0$ then completes the proof of (8).
@@
For the difference in the mutual information, the inequality (1) in Corollary 1.4 states that